Please help me solve this problem:
The smallest mammal, Kitti's hog-nosed bat has a head-body length of 1 7/50inches and a wingspan of about 5 1/10 inches. A scale drawing of this little bat is made showing it in full flight. The wingspan on the drawing is 15 inches. What should the length of the bat be in the drawing?
L= 1 7/50
15= 5 1/10
15 x 1.14 = 17.10 / 5.10 = 3.35 or 3 1/3
is this correct

Answer by stanbon(75887)

Please help me solve this problem:
The smallest mammal, Kitti's hog-nosed bat has a head-body length of
1 7/50inches and a wingspan of about 5 1/10 inches.
--------------------------
A scale drawing of this little bat is made showing it in full flight. The wingspan on the drawing is 15 inches. What should the length of the bat be in the drawing?
---
Use a proportion:
length/length = wing/wing
x/(57/50) = 15/(51/10)
----
x = (57/50)(150/51)
x = (57/1)(3/51)
x = (57/17) = 3.352 inches (length in drawing)

L= 1 7/50
15= 5 1/10
15 x 1.14 = 17.10 / 5.10 = 3.35 or 3 1/3
is this correct
-----
I used a different setup but your answer is correct.
=======
Cheers,
Stan H.
==============